### v2

# Distributed Sign Momentum with Local Steps for Training Transformers

Abstract
Pre-training Transformer models is resource-intensive, and recent studies have shown that sign momentum is an efficient technique for training large-scale deep learning models, particularly Transformers. However, its application in distributed training remains underexplored. This paper investigates a novel communication-efficient distributed sign momentum method with multiple local steps, to cope with the scenarios where communicating at every step is prohibitive. Our proposed method allows for a broad class of base optimizers for local steps, and uses sign momentum in the global step, where momentum is generated from differences accumulated during local steps. For generic base optimizers, by approximating the sign operator with a randomized version that acts as a continuous analog in expectation, we present a general convergence analysis, which specializes to an $O(1/\sqrt{T})$ rate for a particular instance. When local step is stochastic gradient descent, we show an optimal $O(1/T^{1/4})$ rate in terms of $\ell_{1}$ gradient norm for nonconvex smooth cost functions. We extensively evaluate our method on the pre-training of various sized GPT-2 models from scratch, and the empirical results show significant improvement compared to other distributed methods with multiple local steps.
1 Introduction
In this paper, we tackle the distributed optimization problem across $n$ workers:
$\mathop{\mathrm{minimize}}_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x}):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}f_{i}(\bm{x},\bm{\xi}^{(i)}),$ (1)
where $f_{i}$ represents the cost function of a supervised learning model, $\bm{x}$ denotes the model parameters, and $\bm{\xi}^{(i)}$ refers to the data samples drawn from data distribution $\mathcal{D}_{i}$ on worker $i\in[n]$. In distributed training tasks, $\mathcal{D}_{i}$ can be a finite collection of samples allocated on worker $i$.
The problem formulation (1) serves as a natural abstraction for distributed training (Lian et al., 2017; McMahan et al., 2017). A common approach to solving (1) involves workers first computing mini-batch gradients of $f_{i}$, followed by an all-reduce step to aggregate the local mini-batch gradients and obtain the global average. A global optimization step, such as Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951), Momentum (Polyak, 1964), Nesterov’s Accelerated Gradient (NAG) (Nesterov, 1983), Adam (Kingma, 2014) and its weight decay variant AdamW (Loshchilov, 2017), or more recently Lion (Chen et al., 2024b), is then applied. This method is well suited for distributed learning scenarios where communication is inexpensive.
However, in large-scale distributed training across GPU clusters, synchronized communication during the all-reduce step can become a bottleneck due to stragglers (Dean et al., 2012) or slow inter-node and inter-cluster communication, making per-step communication impractical. Several approaches have been proposed to mitigate this communication overhead. Decentralized methods, such as gossiping SGD (Jin et al., 2016) and consensus-based distributed SGD (Jiang et al., 2017), use approximate mini-batch gradient averaging instead of exact averaging to reduce idle time. Local step methods, like local SGD (Lin et al., 2020; Gorbunov et al., 2021b), local momentum (PR-SGD-Momentum) (Yu et al., 2019) and local Adam (Lu et al., 2023), reduce communication frequency to save on bandwidth. Additionally, some local step methods further optimize performance by overlapping communication and computation (Wang et al., 2019; Sun et al., 2024), utilizing stale synchronized model parameters for local steps to reduce the waiting time for synchronizing the latest local models. All these approaches, although reduce communication and run faster, add perturbations to the optimization process. These perturbations can lead to decreased performance in terms of training and validation losses under the same amount of computation (Wang et al., 2020). In this work, we focus on designing local step methods with excellent communication-performance tradeoffs.
Several methods have been proposed to enhance the performance of local steps techniques. SCAFFOLD (Karimireddy et al., 2020) introduces a control variate that corrects the local SGD direction using global information, eliminating the dependency on local gradient dissimilarity in the convergence guarantees. One of the most relevant methods, BMUF (Chen and Huo, 2016), combines local steps with a global momentum step to accelerate distributed training, without the need for additional communication (which can also be a bottleneck in high-dimensional problem), such as transmitting the control variate in SCAFFOLD or the local momentum in PR-SGD-Momentum (Yu et al., 2019). SlowMo (Wang et al., 2019) extends BMUF by providing a general framework that operates on local base optimizers, such as local SGD and SGP (Assran et al., 2019), and periodically applies a global momentum step after running the local optimizer for several iterations.
Beyond local step methods, momentum has long been recognized as a key technique for improving deep learning training, enhancing both optimization and generalization (Sutskever et al., 2013). More recently, Kunstner et al. (2023) suggest that sign momentum may be the key factor behind Adam’s significant advantage over SGD in training Transformers (Vaswani, 2017). Chen et al. (2024b) introduced the Lion (Evolved Sign Momentum) optimization method, discovered through program search, which outperforms the widely used Adam (Kingma, 2014) in various vision and autoregressive language modeling tasks. Furthermore, Kunstner et al. (2023), through experiments with different batch sizes, suggest that Adam’s significant performance advantage over SGD likely stems from its similarity to sign descent with momentum, particularly in Transformers used for language models.
Inspired by the effectiveness of the sign momentum method in training large neural networks, particularly Transformers, and the success of global momentum steps in local step methods, a natural question is:
Can we extend the sign momentum method to distributed training of Transformers with local steps?
1.1 Contributions
In this work, we answer this question affirmatively by proposing a framework for distributed sign momentum with local steps, supported by both empirical evaluations and theoretical justifications. Our contributions are as follows:
  1. (a)
    We introduce a new framework for distributed sign momentum with local steps. This framework is versatile, allowing any proper base optimizers for local steps, while using differences accumulated from these updates to adjust the momentum in the global sign momentum step. Details are provided in Algorithm 1.
  2. (b)
    For generic base optimizer, we provide a general theoretical analysis of the proposed framework by approximating the sign operator with a randomized continuous analog in expectation. When specializing the base optimizer as SGD, we obtain an in-expectation convergence rate of $O\big{(}1/\sqrt{T}\big{)}$ in terms of squared Euclidean norms, where $T$ is the total global steps.
  3. (c)
    When the base optimizer is SGD and the actual sign operator is used, we show an optimal in-expectation rate of $O(1/T^{1/4})$ measured in $\ell_{1}$ gradient norms, matching the best-known rate of centralized sign momentum method.
  4. (d)
    We extensively evaluate the performance of Algorithm 1 in pre-training GPT-2 models of various sizes from scratch, demonstrating consistent improvements over other distributed methods with multiple local steps and highlighting its strength in large-scale scenarios where communicating at every step is prohibitively costly.
1.2 Related work
Communication compression. In distributed optimization, there has been extensive research on compressing communication between workers to reduce overhead. This includes methods such as 1-bit sign SGD (Seide et al., 2014; Bernstein et al., 2018), gradient difference compression (Gorbunov et al., 2021a), compression with error-feedback (Karimireddy et al., 2019), and compression with error reset (Xie et al., 2020), among others. Although this work focuses on local step methods without communication compression, several 1-bit compression (sign-based) methods, such as SIGNUM (Bernstein et al., 2019) and Major Vote signSGD (Sun et al., 2023), apply (random) sign operations to local momentum and use majority voting for local signs. In contrast, Algorithm 1 employs full-precision communication during the global synchronization step, with the sign taken after aggregation.
signSGD and its momentum variants. The convergence of signSGD and its momentum variants has been studied for nonconvex and smooth functions. For signSGD, defined as:
$\bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\nabla f(\bm{x}_{t},\bm{\xi}_{t})),$ (2)
an $O(T^{-1/4})$ convergence rate has been established under an increasing batch size (Bernstein et al., 2018), or under unimodal and symmetric stochastic gradients (Bernstein et al., 2019). For signSGD with a momentum buffer $\bm{m}_{t}$, given by:
$\begin{split}\bm{m}_{t+1}=\beta\bm{m}_{t}+(1-\beta)\nabla f(\bm{x}_{t},\bm{\xi}_{t}),\\ \bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\bm{m}_{t+1}),\end{split}$ (3)
its $O(1/T^{-1/4})$ convergence rate is guaranteed under a decreasing step size and increasing batch size, as shown by Bernstein et al. (2018). Sun et al. (2023) improve the previous results by showing the same convergence rate under a constant batch size independent of $T$. Additionally, Karimireddy et al. (2019) shows that signSGD with error-feedback, another momentum variant, also achieves an $O(T^{-1/4})$ convergence rate, albeit under a more stringent bounded stochastic gradient condition.
Distributed sign momentum. The idea of distributed sign momentum is also studied in works including Liu et al. (2024a) and Tang and Chang (2024). Tang and Chang (2024) uses Lion for multiple local steps and average model and momentum periodically, and Liu et al. (2024a) computes the momentum term in Lion locally for once and uses global aggregation for model updates. Conversely, our proposed framework supports multiple local steps using any suitable base optimizer to generate local differences, which are then leveraged for sign momentum in the global step.
2 Distributed sign momentum
We introduce a general framework, outlined in Algorithm 1, that can incorporate any off-the-shelf base optimizer within the inner loop. Each worker $i\in[n]$ maintains a local model copy $\bm{x}_{t,k}^{(i)}$ for the outer iteration $t$ and inner iteration $k$, denoted as $(t,k)$. Specifically, at $t$-th outer iteration, each worker $i\in[n]$ performs $\tau$ local model updates:
$\bm{x}_{t,k+1}^{(i)}=\bm{x}_{t,k}^{(i)}-\gamma_{t}{\bm{d}}^{(i)}_{t,k},k=0,\ldots,\tau-1,$ (4)
where $\gamma_{t}$ is the local learning rate (LR) and ${\bm{d}}^{(i)}_{t,k}$ is the update direction of local base optimizer. We allow for any base optimizers, such as SGD (with or without momentum), AdamW, and others. For example, if the base optimizer is mini-batch SGD, then
${\bm{d}}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)}),$ (5)
where $\bm{\xi}_{t,k}^{(i)}$ is a sample drawn from $\mathcal{D}_{i}$. As in Wang et al. (2019), base optimizers can also involve communication steps, such as SGP (Assran et al., 2019), and momentum steps. We denote the final update direction applied to worker $i$ at iteration $(t,k)$ as ${\bm{d}}_{t,k}^{(i)}$.
After $\tau$ local steps using the base optimizer, the workers communicate to perform a global sign momentum step. Two global buffers are maintained: the model buffer $\bm{x}_{t,0}$ and the momentum buffer $\bm{m}_{0}$. All workers are assumed to be initialized with $\bm{x}_{0,0}$ and $\bm{m}_{0}=\mathbf{0}$. First, an all-reduce operation is performed to compute the average of the local models, i.e., $\bm{x}_{t,\tau}=(1/n)\sum_{i=1}^{n}\bm{x}_{t,\tau}^{(i)}$. The momentum buffer is then used to execute the global sign momentum step (line 9), for some $\beta_{1}\in[0,1]$,
${\bm{u}}_{t+1}$ $=\beta_{1}\bm{m}_{t}+\frac{1-\beta_{1}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau}),$ (6)
$\bm{x}_{t+1,0}$ $=\bm{x}_{t,0}-\eta\gamma_{t}\big{(}\text{sign}({\bm{u}}_{t+1})+\lambda\bm{x}_{t,0}\big{)},$ (7)
where $\eta$ is the global learning rate and $\lambda$ is the decoupled weight decay. Then, we perform the following global momentum update (line 8), for some $\beta_{2}\in[0,1]$,
$\bm{m}_{t+1}=\beta_{2}\bm{m}_{t}+\frac{1-\beta_{2}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau}).$ (8)
Note that (6)-(8) mimics the update rule of Lion (Chen et al., 2024a) (also described in Section A in Appendix), by treating the differences accumulated from local steps as pseudo-stochastic gradients. In both (6) and (8), we scale $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ by $\frac{1}{\gamma_{t}}$ to make the momentum buffer independent of the learning rate $\gamma_{t}$, which can vary over time when using a learning rate schedule.
Algorithm instances. Instances of Algorithm 1 can be obtained by specifying the parameters $\eta,\gamma_{t},\beta_{1},\beta_{2},\lambda$. For example, setting $\beta_{1}=\beta_{2}=\beta,\lambda=0,\tau=1$ recovers signSGD with momentum as described in (3). In particular, when $n=1$, Algorithm 1 simplifies to a signed Lookahead optimizer (Zhang et al., 2019) with decoupled weight decay.
Algorithm 1 Distributed Sign Momentum
0:  Local update directions ${\bm{d}}_{t,k}^{(i)}$, local learning rate $\gamma$, global learning rate $\eta$, momentum coefficients $\beta_{1},\beta_{2}$.
1:  Initialize $\bm{x}_{0,0}$, let $\bm{m}_{0}=0.$
2:  for $t=0$ to $T-1$ do
3:     for $i=1$ to $n$ in parallel do
4:        for  $k=0$ to $\tau-1$ do
5:           $\bm{x}^{(i)}_{t,k+1}\leftarrow\bm{x}^{(i)}_{t,k}-\gamma_{t}{\bm{d}}_{t,k}^{(i)}$
6:        end for
7:     end for
8:     All-reduce step: $\bm{x}_{t,\tau}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{(i)}_{t,\tau}$.
9:     Update global parameter: $\qquad\bm{u}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+\frac{1-\beta_{1}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ $\qquad\bm{x}_{t+1,0}\leftarrow\bm{x}_{t,0}-\eta\gamma_{t}\big{(}\text{sign}(\bm{u}_{t+1})+\lambda\bm{x}_{t,0}\big{)}$
10:     Update momentum: $\qquad\bm{m}_{t+1}=\beta_{2}\bm{m}_{t}+\frac{1-\beta_{2}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$
11:     Synchronize over all workers: $\qquad\forall i\in[n],\ \bm{x}^{(i)}_{t+1,0}\leftarrow\bm{x}_{t+1,0}$
12:  end for
13:  return $\bm{x}_{T,0}$
Global sign momentum step. The updates (6)-(8) offer a flexible configuration for the global sign momentum step. Specifically, when $\beta_{2}>\beta_{1}$, the current $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ takes on a larger weight in the exponential moving average, similar to classical momentum as in SlowMo. This adjustment in momentum can lead to further acceleration, as studied by Shi et al. (2022). Additionally, the decoupled weight decay in (7) is similar to the weight decay used in AdamW, making it more advantageous than Adam.
Distributed optimizer. We note that Algorithm 1 can be seamlessly integrated with the ZERO-series distributed optimizer (Rajbhandari et al., 2020) to handle large models. Consider a layered distributed optimization strategy: GPUs on a single node can be treated as a joint worker, utilizing a distributed optimizer like ZERO2 for local steps of the base optimizer, while the global buffers $\bm{x}_{t,0}$ and $\bm{m}_{t,0}$ are distributed across nodes. This approach allows more frequent communication for local steps to leverage faster intra-node communication, such as NVLinks.
3 Convergence analysis
We provide some theoretical justifications on the convergence of Algorithm 1 under some regular conditions. Let $f_{i}(\bm{x}):=\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}f_{i}(\bm{x},\bm{\xi}^{(i)})$. We make the following assumptions on local functions.
Assumption 1.
Each local function $f_{i}$ is differentiable and $\forall\bm{x},\bm{y}\in\mathbb{R}^{d},\|\nabla f_{i}(\bm{x})-\nabla f_{i}(\bm{y})\|\leq L\|\bm{x}-\bm{y}\|$ for some $L>0$.
Let $\mathbb{E}_{t,k}$ denote the conditional expectation given all the historical randomness up to iteration $(t,k)$. We define the virtual global averages:
$\bm{x}_{t,k}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{t,k}^{(i)},\ {\bm{d}}_{t,k}=\frac{1}{n}\sum_{i=1}^{n}{\bm{d}}_{t,k}^{(i)}.$
We assume that the average update direction ${\bm{d}}_{t,k}$ has bounded variance.
Assumption 2.
There exists some $\zeta>0$ such that $\mathbb{E}[\|{\bm{d}}_{t,k}-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}]\leq\zeta^{2}$.
Similar to Sun et al. (2023), we assume that update directions of base optimizers are uniformly bounded throughout the optimization process.
Assumption 3.
The local update direction of each local base optimizer is bounded at at all iteration $(t,k)$: $\forall(t,k),\forall i\in[n],\ \|{\bm{d}}_{t,k}^{(i)}\|^{2}\leq R^{2}$ for some $R>0$.
3.1 Randomized sign operator
For a generic base optimizer that generates update directions satisfying Assumption 2 and 3, we present an general analysis that approximates the sign operator with a randomized analog. For any vector ${\bm{v}}=[v_{1},\ldots,v_{d}]^{\top}\in\mathbb{R}^{d}$ such that $\|{\bm{v}}\|\leq B$, we use the randomized sign operator $\mathcal{S}_{r}({\bm{v}})$, where the $j$-th component is defined as:
$[\mathcal{S}_{r}({\bm{v}})]_{j}=\begin{cases}-\mathrm{sign}(v_{j}),&\text{with probability (w.p.)}\frac{1}{2}-\frac{|v_{j}|}{2B},\\ \mathrm{sign}(v_{j}),&\text{ w.p. }\frac{1}{2}+\frac{|v_{j}|}{2B}.\end{cases}$ (9)
Or alternatively,
$[\mathcal{S}_{r}({\bm{v}})]_{j}=\begin{cases}0,&\text{ w.p. }1-\frac{|v_{j}|}{B},\\ \mathrm{sign}(v_{j}),&\text{ w.p. }\frac{|v_{j}|}{B}.\end{cases}$ (10)
We use these randomized sign operators as a linear and continuous analog of the original sign operator in expectation as in Safaryan and Richtárik (2021). Let $\mathbb{E}_{\mathcal{S}}$ denote the expectation taken over the randomness of the randomized sign operator, we have the following property.
Lemma 1.
For a random vector $\bm{v}\in\mathbb{R}^{d}$ that satisfies $\|{\bm{v}}\|\leq B$ almost surely, we have $\mathbb{E}_{\mathcal{S}}[{\bm{v}}]={\bm{v}}/B$, and $\mathbb{E}_{\mathcal{S}}[\|\mathcal{S}_{r}({\bm{v}})-{\bm{v}}/B\|^{2}]\leq d$.
With the above linerization in expectation, we establish a general convergence for some specific instances of Algorithm 1. Proofs of all theorems are deferred to the Appendix.
Theorem 1.
Let Assumptions 1, 2, and 3 hold. Run Algorithm 1 by approximating sign with $\mathcal{S}_{r}$ in (9) or (10) with $B=\tau R$. Take local learning rate $\gamma=\frac{R}{\eta}\sqrt{\frac{n\tau}{T}}$, weight decay $\lambda=0$, and momentum coefficients $\beta_{1}=\beta_{2}=\beta\in[0,1)$. When the outer iterations number $T\geq 4nL^{2}\big{[}4(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{8\tau\beta^{2}}{(1-\beta)^{2}}+1\big{]}$, the iterates $\{\bm{x}_{t,k}\}$ generated from Algorithm 1 satisfy that
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}\leq\frac{2(f(\bm{x}_{0,0})-f_{*})}{\sqrt{n\tau T}}$
$\quad+\underbrace{\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]}_{\text{Effect of base optimizer}}+\zeta^{2}L\sqrt{\frac{n}{\tau T}}$
$\quad+\frac{4nL^{2}\zeta^{2}}{T}\Big{[}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{2\beta^{2}}{1-\beta^{2}}\Big{]}+dLR^{2}\sqrt{\frac{n\tau}{T}}.$
Theorem 1 provides a general convergence guarantee for the simple sign momentum case of Algorithm 1. The convergence rate depends on the effect of base optimizer, which can measure both gradient heterogeneity across workers and stochastic gradient bias. We next specialize to the case where the base optimizer is SGD.
Table 1: Model configurations and peak learning rates
Size Embedding dimension # of heads Depth Peak LR
Small 125M 768 12 12 5e-4
Medium 355M 1024 16 24 2e-4
Large 770M 1280 20 36 2e-4
Theorem 2.
Let Assumptions 1, 3 hold, and in addition assume that:
  1. (a)
    the local update direction ${\bm{d}}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$, and $\forall\bm{x}\in\mathbb{R}^{d},\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}[\|\nabla f_{i}(\bm{x},\bm{\xi}^{(i)})-\nabla f_{i}(\bm{x})\|^{2}]\leq\sigma^{2}$ for some $\sigma>0$;
  2. (b)
    $\forall\bm{x}\in\mathbb{R}^{d},(1/n)\sum_{i=1}^{n}\|\nabla f(\bm{x})-\nabla f_{i}(\bm{x})\|^{2}\leq\delta^{2}$ for some $\delta>0$.
Then, by approximating sign with randomized operator $\mathcal{S}_{r}$ and using parameters $\gamma,\eta,\lambda,\beta_{1},\beta_{2}$ and $T$ as in Theorem 1, we have $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})\|^{2}]=O(1/\sqrt{T})$, specifically, it
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}\leq\frac{2(f(\bm{x}_{0,0})-f_{*})}{\sqrt{n\tau T}}$
$\quad+\frac{3n\tau^{2}L^{2}R^{2}(\sigma^{2}+3\tau\delta^{2})}{\eta^{2}T}+\sigma^{2}L\sqrt{\frac{n}{\tau T}}$
$\quad+\frac{4nL^{2}\sigma^{2}}{T}\Big{[}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{2\beta^{2}}{1-\beta^{2}}\Big{]}+dLR^{2}\sqrt{\frac{n\tau}{T}}.$
Remark 1.
We compare the instance in Theorem 2 with a related algorithm, Federated Majority Vote signSGD with Stochastic Simple Momentum (Federated MV-sto-signSGD-SIM), analyzed by Sun et al. (2023). Federated MV-sto-signSGD-SIM incorporates similar algorithmic concepts but remains distinct from Algorithm 1. First, our algorithm uses aggregated local differences $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ for a global momentum update while MV-sto-signSGD-SIM uses local stochastic gradients to update local momentum. Second, Sun et al. (2023) uses randomized sign to compress worker to server communications, i.e., sign based majority vote, while we use full-precision communication among workers.
3.2 Sign operator
When the base optimizer is SGD, we also present its convergence analysis using actual sign for Algorithm 1.
Theorem 3.
Let Assumptions 1, 3 hold. In addition, the local update on worker $i$ satisfies that ${\bm{d}}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$, and $\forall\bm{x}\in\mathbb{R}^{d},\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}[\|\nabla f_{i}(\bm{x},\bm{\xi}^{(i)})-\nabla f_{i}(\bm{x})\|^{2}]\leq\sigma^{2}$ for some $\sigma>0$. Taking $\eta=\frac{1}{LT^{3/4}}$, and $1-\beta=\frac{1}{\sqrt{T}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{L(f(\bm{x}_{0,0})-f_{*})}{\gamma T^{1/4}}$
$\quad+\frac{2\sqrt{d}}{T^{1/2}}\|\nabla f(\bm{x}_{0,0})\|+\frac{2d\gamma}{T^{1/4}}$
$\quad+\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}+\frac{\sqrt{d}\tau R+\gamma d/2}{T^{3/4}}.$
Remark 2.
The convergence rate above in terms of $\ell_{1}$ gradient norm matches the rate of centralized sign momentum method, namely, update in (3), as shown in Sun et al. (2023). Specifically, in one of the leading terms $\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}$, the benefit of multiple local steps and linear speedup are demonstrated, and it suggests our algorithm is preferable in large noise regime. We note that Sun et al. (2023) also shows similar convergence rate of federated MV-sto-signSGD-SIM), but only to an $O(\frac{dR}{\sqrt{n}})$ neighborhood under the same Assumption 3, due to its use of randomized signs for worker to server communication.
4 Experiments
We evaluate Algorithm 1 on auto-regressive language modeling using GPT-2 (Radford et al., 2019), with model sizes ranging from 125M to 770M parameters, trained from scratch on the OpenWebText dataset (Gokaslan et al., 2019). Following standard practice, we set the context length of GPT-2 to 1024 tokens. We consider three versions of GPT-2: 125M (Small), 355M (Medium), and 770M (Large) parameters, with configurations as detailed in Table 1.
Figure 1: Validation loss curves versus communication rounds for GPT-2 small, medium, and large. Communication interval for our Algorithm 1 and SlowMo are set as $\tau=12$.
Figure 2: Validation loss curves versus computation rounds for GPT-2 small, medium, and large. Communication interval for our Algorithm 1 and SlowMo are set as $\tau=12$.
Baselines. We benchmark the empirical performance of our proposed framework, namely, Algorithm 1, using AdamW as the base optimizer, against the state-of-the-art distributed training method with multiple local steps, i.e., SlowMo (Algorithm 5 in Appendix) that also uses AdamW as the base optimizer. AdamW is the dominant optimizer for the pre-training of Transformers (Liu et al., 2024b), and according to (Sun et al., 2024), SlowMo with AdamW as the base optimizer is among the best performing distributed training methods with multiple local steps. We measure practical performance of Algorithm 1 and SlowMo in terms of validation losses, i.e., token-level log perplexity. In general, under the same compute budget, due to significant communication savings, distributed methods with multiple local steps suffer from performance drop compared to those methods that communicate in every computation round. To demonstrate that our proposed Algorithm 1 suffers from much less performance drop, we also report the validation losses of running standalone AdamW optimizer that communicates (all-reduce mini-batch gradients) in every computation round. We omit the comparison with local AdamW, i.e., averaging models over workers after every $\tau$ AdamW steps at each worker, due to its poor performance.
Alg. Com. red. GPT-2 Small 124M GPT-2 Medium 355M GPT-2 Large 770M
Val. Improv. Val. Improv. Val. Improv.
AdamW N.A. 2.917 2.682 2.552
  SlowMo 12×12$\times$12 × 2.993 2.810 2.727
Algorithm 1 12×12$\times$12 × 2.942 5.23% 2.709 10.63% 2.658 7.14%
  SlowMo 24×24$\times$24 × 2.995 2.800 2.733
Algorithm 1 24×24$\times$24 × 2.957 3.8% 2.725 7.79% 2.864 14.00%
  SlowMo 36×36$\times$36 × 3.000 2.818 2.736
Algorithm 1 36×36$\times$36 × 2.972 2.84% 2.750 7.04% 2.714 2.22%
Table 2: Performance comparison of standalone AdamW (with per iteration communcation), SlowMo + AdamW, and Algorithm 1 + AadmW under communcation intervals $\tau=12,24,36$ for GPT-2 small, medium, and large. Com. red. means communication reduction, and Val. denotes final validation losses, Improv. is the improvement of Algorithm 1 over baseline SlowMo in terms of perplexity, i.e., the natural exponential of validation loss improvement.
Implementations. We set the total batch size to 480 for all algorithms, and run each for 100k steps to ensure the same compute budget, as well as the same total number of tokens. In this setup, we fine-tune the recommended parameters for AdamW (both for the base optimizer AdamW in Algorithm 1 and SlowMo): $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a cosine learning rate schedule (Loshchilov and Hutter, 2016) with a 2k-step warm-up. The peak learning rate (LR) is specified in Table 1, and the final LR is set to 0.05×0.05$\times$0.05 × the peak LR. For the global sign momentum step in Algorithm 1, we apply the recommended Lion parameters: $\beta_{1}=0.95$, $\beta_{2}=0.98$, and $\lambda=0.1$ (Liu et al., 2024b). We tune the momentum coefficient and global LR for SlowMo, and the global LR for Algorithm 1. Performance is tested on three different sizes of GPT-2 models across varying numbers of workers.
Parameter tuning. We performn the following model pre-training.
  1. 1.
[…]
  2. 2.
    GPT-2 Medium on 8 workers. For SlowMo, we use the best performing momentum coefficient in GPT-2 Small training, and tune global LR in $\{0.5,0.8,1.0\}$. For Algorithm 1, we tune global LR in $\{0.5,0.8,1.0,1.2,1.5,2.0\}$.
  3. 3.
    GPT-2 Large on 16 workers. SlowMo momentum coefficient is chosen from $\{0.5,0.6\}$ and global LR is from $\{0.5,1.0\}$. Algorithm 1 uses the best global LR in $\{0.5,0.8,1.0,1.2\}$.
Performance improvements. In Figure 1 and 2, we present the validation loss curves against communication rounds and computation rounds, respectively, during the pre-training of three GPT-2 models. The communication interval $\tau$ for both our proposed Algorithm 1 and SlowMo is set to 12. Figure 1 and 2 show that our proposed Algorithm 1 saves 12×12$\times$12 × communications compared to standalone AdamW that communicates in every computation round, and achieves comparable performance. In all three experiments, our method consistently outperforms SlowMo by a significant margin, and suffers less than half performance drop compared to the standalone AdamW, particularly for GPT-2 Small and GPT-2 Medium. For example, for GPT-2 medium, Algorithm 1 achieves final validation loss 2.709, while SlowMo only reaches 2.810. The 0.1 validation loss improvement is significant, since according to scaling laws Liu et al. (2024b); Hoffmann et al. (2022), even an 0.05 validation loss improvement can take 2×2$\times$2 × compute to achieve.
On the effects of communication interval $\tau$. In Table 2, we report the final validation losses of training three sized GPT-2 models with various algorithm setups. It shows that Algorithm 1 achieves consistent and robust performance improvements versus SlowMo under different communication intervals, $\tau=12,24,36$, across all model sizes, highlighting the advantages of sign momentum over standard momentum. Additionally, we observe that as $\tau$ increases, the performance gap between our method and SlowMo narrows. Nevertheless, a 10×10$\times$10 × or 20×20$\times$20 × reduction in communication would still offer substantial communication savings in practical distributed training scenarios where per-iteration communication is prohibitively expensive.
More base optimizers. We also test our proposed framework Algorithm 1 using base optimizer Sophia (Liu et al., 2024b), a more recent optimizer shown to outperform AdamW. We train a GPT-2 small model over 4 workers, and compare with SlowMo that also uses Sophia base optimizer under communication interval $\tau=12$, and standalone Sophia with per-iteration mini-batch gradient all-reduce. We use the suggested hyper parameters of Sophia, except tuning its peak LR and accordingly the minimum LR. Similar to previous experiments, we tune momentum coefficient and global LR for SlowMo and global LR for Algorithm 1. We report validation losses in Table 3. Although both Algorithm 1 and SlowMo achieve worse losses than the case of using AdamW as the base optimizer, Algorithm 1 still improves over SlowMo over 5%.
Alg. Com. red. Val. Improv.
Sophia N.A. 2.873
SlowMo 12×12$\times$12 × 3.016
Algorithm 1 12×12$\times$12 × 2.965 5.23%
Table 3: Performance comparison of SlowMo and Algorithm 1 with Sophia as the base optimizer.
4.1 Ablation studies
We conduct several ablation studies to explore the factors contributing to the effectiveness of sign momentum. In all these studies, we used the AdamW optimizer as the base optimizer, as described in the previous section.
Lookahead and signed Lookahead optimizer. We begin by testing whether momentum and signed momentum improve methods with multiple local steps in the case of a single worker. To do this, we configure Algorithm 1 with $n=1$, $\beta_{1}=\beta_{2}=\beta$, and $\lambda=0$, reducing it to the signed Lookahead optimizer. Upon that, by replacing (7) with $\bm{x}_{t+1,0}=\bm{x}_{t,0}-\eta\gamma_{t}{\bm{u}}_{t+1}$, we recover the Lookahead optimizer (Zhang et al., 2019). We test the Lookahead optimizer using base optimizer AdamW in training GPT-2 medium, under $\tau=48$, global LR = 1, with validation losses reported in Table 4. We also evaluate the signed Loookahead optimizer with AdamW as the base optimizer in training a GPT-2 small model with $\tau=24$, global LR = 6, and the validations losses are reported in Table 5. These results show that both Lookahead and signed Lookahead are effective in improving the performance of base optimizer when $n=1$, highlighting the critical role of momentum in training Transformers.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.682
Lookahead $0.1$ 2.665 1.71%
Lookahead $0.2$ 2.660 2.22%
Table 4: Performance comparison of AdamW and Lookahead with AadmW as the base optimizer.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.917
Signed Lookahead $0.6$ 2.896 2.12%
Signed Lookahead $0.8$ 2.890 2.74%
Table 5: Performance comparison of AdamW and signed Lookahead with AadmW as the base optimizer.
Signed SlowMo. We now consider the case with multiple workers, i.e., when $n>1$. In this setting, we define signed SlowMo by setting $\beta_{1}=\beta_{2}=\beta$, and modifying the updates in (6) and (7) as follows:
${\bm{u}}_{t+1}$ $=\beta_{1}\bm{m}_{t}+\frac{1-\beta_{1}}{\gamma_{t}}\text{sign}(\bm{x}_{t,0}-\bm{x}_{t,\tau}),$
$\bm{x}_{t+1,0}$ $=\bm{x}_{t,0}-\eta\gamma_{t}{\bm{u}}_{t+1}.$
We use signed SlowMo to pre-train the GPT-2 Small model with a communication interval of $\tau=12$, and report final validation losses in Table 6. It shows that signed SlowMo achieves better final validation losses compared to SlowMo, showing the efficacy of incorporating sign momentum. However, it still underperforms compared to our proposed Algorithm 1 (best loss 2.942), which we hypothesize is due to the acceleration effects provided by using $\beta_{2}>\beta_{1}$.
Alg. $\beta$ Val. Improv.
AadmW N.A. 2.917
SlowMo 2.993
Signed SlowMo 0.5 2.974 1.92%
Signed SlowMo 0.8 2.970 2.32%
Global AdamW N.A 2.992 0.10%
Table 6: Performance comparison of AdamW; SlowMo, signed SlowMo, and global AdamW with AadmW as the base optimizer.
Adaptive global update. We further compare the global sign momentum method with a global step that mimics AdamW (see Appendix for a description). Balles and Hennig (2018) show that Adam can be regarded as sign momentum with adaptive learning rate depending on relative variance. Its performance, as reported in Table 6, is only comparable with SlowMo, partially indicating that this adaptivity brings limited benefits when used as the global step of local step methods. We defer its detailed description of this comparisons to the Appendix.
5 Conclusion and Discussions
In this paper, we have introduced a new framework for distributed sign momentum with local steps, integrating sign momentum with a communication-efficient local step approach. The framework is flexible, accommodating a variety of off-the-shelf base optimizers. Comprehensive empirical results from pre-training GPT-2 models of various sizes from scratch consistently show improvements over the baseline method, validating the effectiveness of our approach in scenarios where communicating at every step is impractical. Additionally, a series of ablation studies highlights the impact of sign momentum in different scenarios. We have provided a general theoretical analysis for generic base optimizers, by utilizing a randomized version of the sign operator, and provided an $O(1/\sqrt{T})$ convergence rate when using SGD as the base optimizer. For SGD as base optimizer in particular, under actual sign operator, we have shown an optimal $O(1/T^{1/4})$ convergence that matches the rate of centralized method. To the best of our knowledge, this framework is a novel contribution to the literature.
Future directions include evaluating the performance of the proposed method on training vision models, such as Vision Transformers (Dosovitskiy, 2020), and investigating the convergence of Algorithm 1 under broader parameter settings and with the real sign operator.
Appendix A Optimizer descriptions
In this section, we provide detailed descriptions of the algorithms referenced in the main text. In the absence of local steps, we present the pseudocode for Polyak’s Momentum (Algorithm 3), AdamW (Algorithm 2), and Lion (Algorithm 4). We denote the stochastic gradient computed at $\bm{x}$ using random samples $\bm{\xi}$ as $\nabla f(\bm{x},\bm{\xi})$, where $\bm{x}$ and $\bm{\xi}$ are defined based on the given context. For distributed algorithms with local steps, we give the pseducode of SlowMo (Algorithm 5) and Federated Majority Vote signSGD with Stochastic Simple Momentum (Federated MV-sto-signSGD-SIM in Algorithm 6).
Algorithm 2 AdamW, Adam (Kingma, 2014) with decoupled weight decay proposed by Loshchilov (2017). Component-wise vector multiplication $\bm{g}_{t}^{2}=\bm{g}_{t}\odot\bm{g}_{t}$, and $\sqrt{\widehat{\bm{v}}_{t}}$ means component-wise square root.
0:  Initialization $\bm{x}_{0}$, momentum coefficients $\beta_{1},\beta_{2}$; weight decay $\lambda$; learning rate $\eta$; $\epsilon=10^{-8}$
  $\bm{m}_{0}\leftarrow\mathbf{0},\bm{v}_{0}\leftarrow\mathbf{0}$
  for $t=0$ to $T-1$ do
     Stochastic gradient: $\bm{g}_{t}\leftarrow\nabla f(\bm{x}_{t},\bm{\xi}_{t})$
     Momentum updates: $\qquad\bm{m}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+(1-\beta_{1})\bm{g}_{t}$ $\qquad\bm{v}_{t+1}\leftarrow\beta_{2}\bm{v}_{t}+(1-\beta_{2})\bm{g}_{t}^{2}$
     Bias-corrections: $\qquad\widehat{\bm{m}}_{t+1}\leftarrow\bm{m}_{t+1}/(1-\beta_{1}^{t+1})$ $\qquad\widehat{\bm{v}}_{t+1}\leftarrow\bm{v}_{t+1}/(1-\beta_{2}^{t+1})$
     Parameter update: $\qquad\bm{x}_{t+1}\leftarrow\bm{x}_{t}-\eta\Big{(}\frac{\widehat{\bm{m}}_{t+1}}{\sqrt{\widehat{\bm{v}}_{t+1}}+\epsilon}+\lambda\bm{x}_{t}\Big{)}$
  end for
  return $\bm{x}_{T}$
0:  Initialization $\bm{x}_{0}$, momentum coefficient $\beta$, learning rate $\eta>0$
1:  $\bm{m}_{0}\leftarrow\mathbf{0}$
2:  for $t=0$ to $T-1$ do
3:     Stochastic gradient: $\bm{g}_{t}\leftarrow\nabla f(\bm{x}_{t},\bm{\xi}_{t})$
4:     Momentum update: $\bm{m}_{t+1}\leftarrow\beta\bm{m}_{t}+\bm{g}_{t}$
5:     Parameter update: $\bm{x}_{t+1}\leftarrow\bm{x}_{t}-\eta\bm{m}_{t+1}$
6:  end for
7:  Return $\bm{x}_{T}$
Algorithm 3 Polyak’s Momentum (Polyak, 1964).
Algorithm 4 Lion (Chen et al., 2024b).
0:  Initialization $\bm{x}_{0}$; momentum coefficients $\beta_{1},\beta_{2}$; weight decay $\lambda$, learning rate $\eta$
1:  $\bm{m}_{0}\leftarrow\mathbf{0}$
2:  for $t=0$ to $T-1$ do
3:     Stochastic gradient: $\bm{g}_{t}\leftarrow\nabla f(\bm{x}_{t},\bm{\xi}_{t})$
4:     Parameter update: $\qquad{\bm{u}}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+(1-\beta_{1})\bm{g}_{t}$ $\qquad\bm{x}_{t+1}\leftarrow\bm{x}_{t}-\eta(\text{sign}(\bm{u}_{t+1})+\lambda\bm{x}_{t})$
5:     Momentum update: $\qquad\bm{m}_{t+1}\leftarrow\beta_{2}\bm{m}_{t}+(1-\beta_{2})\bm{g}_{t}$
6:  end for
7:  Return $\bm{x}_{T}$
Algorithm 5 SlowMo, a general framework that combines local steps and global momentum step (Wang et al., 2019). Worker $i$’s updating direction of base optimizer at iteration $(t,k)$ is denoted as ${\bm{d}}_{t,i}^{(i)}$.
0:  Initialization $\bm{x}_{0,0}$; base optimizer learning rate $\gamma_{t}$; inner loop steps $\tau$; global learning rate $\alpha$; initial momentum $\bm{u}_{0}=\mathbf{0}$
  for $t=0$ to $T-1$ at worker $i$ in parallel do
     for $k=0$ to $\tau-1$ do
        Base optimizer step: $\bm{x}_{t,k+1}^{(i)}\leftarrow\bm{x}_{t,k}^{(i)}-\gamma_{t}{\bm{d}}_{t,k}^{(i)}$
     end for
     Exact averaging: $\bm{x}_{t,\tau}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{t,\tau}^{(i)}.$
     Update slow momentum: $\qquad\bm{u}_{t+1}\leftarrow\beta\bm{u}_{t}+\frac{1}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$
     Update outer iterate: $\bm{x}_{t+1,0}\leftarrow\bm{x}_{t,0}-\alpha\gamma_{t}\bm{u}_{t+1}$
  end for
  return $\bm{x}_{T,0}$
0:  Iterate initialization $\bm{x}_{-1}=\bm{x}_{0}=\mathbf{0}$, momentum initialization $\bm{m}_{0}=\mathbf{0}$, momentum coefficient $\beta\in[0,1)$, local learning rate $\gamma$, inner loop steps $\tau$, outer momentum coefficient $\alpha$, global learning rate $\eta$, the uniform $\ell_{2}$ norm bound on stochastic gradient $B$
1:  for $t=0$ to $T-1$ do
2:     $\bm{y}_{t}\leftarrow\bm{x}_{t}+\alpha(\bm{x}_{t}-\bm{x}_{t-1})$
3:     for worker $i=1$ to $n$ in parallel do
4:        ${\bm{z}}_{t,0}^{(i)}\leftarrow\bm{y}_{t}$
5:        for $k=0$ to $\tau-1$ do
6:           $\bm{z}_{t,k+1}^{(i)}\leftarrow\bm{z}_{t,k}^{(i)}-\gamma\nabla f_{i}(\bm{z}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$
7:        end for
8:        $\bm{y}_{t}^{(i)}\leftarrow\bm{z}_{t,\tau}^{(i)}$
9:        $\bm{m}_{t+1}^{(i)}\leftarrow\beta\bm{m}_{t}^{(i)}+(1-\beta)\nabla f_{i}(\bm{y}_{t}^{(i)},\bm{\xi}_{t}^{(i)})$
10:     end for
[…]
12:  end for
13:  Return $\bm{x}_{T}$
Algorithm 6 Federated MV-sto-signSGD-SIM (Sun et al., 2023). The randomized sign operator $\mathcal{S}_{r}$ is as defined in (9).
Appendix B Additional proofs
B.1 Proof of Theorem 1
B.1.1 Proof of Lemma 1
Proof.
For $\mathcal{S}_{r}$ defined in (9) and (10), it can be readily verified that if $\|\bm{v}\|\leq B$, $\mathbb{E}_{\mathcal{S}}[\mathcal{S}(\bm{v})]=\bm{v}/B$. Then,
$\mathbb{E}_{\mathcal{S}}[\|\mathcal{S}_{r}(\bm{v})-\frac{\bm{v}}{B}\|^{2}]=\mathbb{E}_{\mathcal{S}}[\|\mathcal{S}_{r}(\bm{v})-\mathbb{E}_{\mathcal{S}}[\mathcal{S}_{r}(\bm{v})]\|^{2}]=\mathbb{E}_{\mathcal{S}}[\|\mathcal{S}_{r}(\bm{v})\|^{2}]-\frac{\|\bm{v}\|^{2}}{B^{2}}\leq d,$
where the last step follows from that $\mathcal{S}_{r}(\bm{v})$’s component is either 0 or ±$\pm 1$. ∎
B.1.2 Virtual iterates
Recall that at iteration $(t,k)$, the local updates are given by $\bm{x}_{t,k+1}^{(i)}=\bm{x}_{t,k}^{(i)}-\gamma{\bm{d}}_{t,k}^{(i)}$. We define the global averages:
${\bm{d}}_{t,k}:=\frac{1}{n}\sum_{i=1}^{n}{\bm{d}}_{t,k}^{(i)},\ \bm{x}_{t,k}:=\frac{1}{n}\sum_{i=1}^{n}{\bm{x}}_{t,k}^{(i)}.$
Then, one has
$\bm{x}_{t,k}=\bm{x}_{t,0}-\gamma\sum_{i=0}^{k-1}\bm{d}_{t,i}.$
In the global update step,
$\bm{x}_{t+1,0}=\bm{x}_{t,0}-\eta\gamma\mathcal{S}_{r}(\bm{m}_{t+1})=\bm{x}_{t,0}-\eta\gamma\mathcal{S}_{r}\Big{(}\beta\bm{m}_{t}+(1-\beta)\sum_{i=0}^{\tau-1}{\bm{d}}_{t,i}\Big{)}.$
By Assumption 3 and Jensen’s inequality, we have $\|{\bm{d}}_{t,i}\|\leq R$ almost surely (a.s.), and thus $\|\sum_{i=0}^{\tau-1}{\bm{d}}_{t,i}\|\leq\tau R$ a.s.. Further, $\bm{m}_{t+1}$ is a convex combination of $\bm{m}_{t}$ and $\sum_{i=0}^{\tau-1}{\bm{d}}_{t,i}$ with $\bm{m}_{0}=\mathbf{0}$, then one can readily show that $\|\bm{m}_{t}\|\leq\tau R$ for all $t\geq 0$. Using $\mathcal{S}_{r}$ with $B=\tau R$ and taking total expectation, we have
$\begin{split}\mathbb{E}[\bm{x}_{t+1,0}]&=\mathbb{E}[\bm{x}_{t,0}]-\eta\gamma\mathbb{E}[\mathcal{S}_{r}(\bm{m}_{t+1})]\\ &=\mathbb{E}[\bm{x}_{t,0}]-\frac{\eta\gamma}{\tau R}\mathbb{E}[\bm{m}_{t+1}]\\ &=\mathbb{E}[\bm{x}_{t,0}]-\frac{\eta\gamma}{\tau R}\Big{(}\beta\mathbb{E}[\bm{m}_{t}]+(1-\beta)\sum_{i=0}^{\tau-1}\mathbb{E}[{\bm{d}}_{t,i}]\Big{)}\end{split}$ (11)
We define the following auxiliary iterates $\bm{y}_{t,k}$ at iteration $(t,k)$ for $k=0,\ldots,\tau$:
$\bm{y}_{t,k}:=\bm{x}_{t,0}-\frac{\eta\gamma}{\tau R(1-\beta)}\big{[}\beta\bm{m}_{t}+(1-\beta)\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\big{]},$ (12)
where we use the convention $\sum_{i=0}^{-1}{\bm{d}}_{t,i}=\mathbf{0}$. From the definition in (12), we have the following relations for the virtual iterates $\{\bm{y}_{t,k}\}$:
$\bm{y}_{t,k+1}=\bm{y}_{t,k}-\frac{\eta\gamma}{\tau R}{\bm{d}}_{t,k},\forall k=0,\ldots,\tau-1,.$ (13)
In addition, under total expectation one has,
[…]
B.1.3 Decent Lemma
Let $\mathbb{E}_{t,k}$ denote the conditional expectation given history up to time $(t,k)$, and $\mathbb{E}_{\mathcal{S}}$ denote the expectation over the random sign operator $\mathcal{S}_{r}$. Since each $f_{i}$ is $L$-smooth, the function $f$ is also $L$-smooth. We have for $k=0,\ldots,\tau-1$,
$\mathbb{E}_{t,k}\big{[}f(\bm{y}_{t,k+1})]-f(\bm{y}_{t,k})$ $\leq-\frac{\eta\gamma}{\tau R}\langle\nabla f(\bm{y}_{t,k}),\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\rangle+\frac{L\eta^{2}\gamma^{2}}{2\tau^{2}R^{2}}\mathbb{E}_{t,k}\big{[}\|{\bm{d}}_{t,k}\|^{2}].$ (15)
For the first term on the right hand side, we have
$\begin{split}&-\langle\nabla f(\bm{y}_{t,k}),\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\rangle\\ =\ &-\langle\nabla f(\bm{y}_{t,k})-\nabla f(\bm{x}_{t,k}),\mathbb{E}_{t,k}[\bm{d}_{t,k}]\rangle-\langle\nabla f(\bm{x}_{t,k}),\mathbb{E}_{t,k}[\bm{d}_{t,k}]\rangle\\ \leq\ &\|\nabla f(\bm{y}_{t,k})-\nabla f(\bm{x}_{t,k})\|^{2}-\frac{1}{4}\|\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2}-\frac{1}{2}\|\nabla f(\bm{x}_{t,k})\|^{2}+\frac{1}{2}\|\nabla f(\bm{x}_{t,k})-\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2}\\ \leq\ &L^{2}\|\bm{y}_{t,k}-\bm{x}_{t,k}\|^{2}-\frac{1}{4}\|\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2}-\frac{1}{2}\|\nabla f(\bm{x}_{t,k})\|^{2}+\frac{1}{2}\|\nabla f(\bm{x}_{t,k})-\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2},\end{split}$ (16)
where the last inequality also follows from the $L$-smoothness of $f$. We first have
$\begin{split}\|\bm{y}_{t,k}-\bm{x}_{t,k}\|^{2}&=\|\gamma\big{(}1-\frac{\eta}{\tau R}\big{)}\sum_{i=0}^{k-1}\bm{d}_{t,i}-\frac{\eta\gamma\beta}{\tau R(1-\beta)}\bm{m}_{t}\|^{2}\\ &\leq 2\gamma^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}\|\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\|^{2}+\frac{2\eta^{2}\gamma^{2}\beta^{2}}{\tau^{2}R^{2}(1-\beta)^{2}}\|\bm{m}_{t}\|^{2}.\end{split}$ (17)
We next bound the second term in (15).
$\begin{split}\mathbb{E}_{t,k}[\|\bm{d}_{t,k}\|^{2}]=\|\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2}+\mathbb{E}_{t,k}\big{[}\|\bm{d}_{t,k}-\mathbb{E}_{t,k}[\bm{d}_{t,k}]\|^{2}\big{]}\end{split}$ (18)
Let $\alpha:=\frac{\eta\gamma}{\tau R}$. Combing (15)-(18) leads to that
$\mathbb{E}_{t,k}\big{[}f(\bm{y}_{t,k+1})-f(\bm{y}_{t,k})\big{]}$ $\leq-\frac{\alpha}{2}\|\nabla f(\bm{x}_{t,k}|\|^{2}-\frac{\alpha}{2}\big{(}\frac{1}{2}-\alpha L\big{)}\|\mathbb{E}_{t,k}\big{[}{\bm{d}}_{t,k}\big{]}\|^{2}+\frac{\alpha^{2}L}{2}\mathbb{E}_{t,k}\big{[}\|{\bm{d}}_{t,k}-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}$
$\quad+\frac{\alpha}{2}\|\nabla f(\bm{x}_{t,k})-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}$
$\quad+2\alpha\gamma^{2}L^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}\|\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\|^{2}+\frac{2\alpha^{3}\beta^{2}L^{2}}{(1-\beta)^{2}}\|\bm{m}_{t}\|^{2}.$
Taking the total expectation,
$\begin{split}\mathbb{E}[f(\bm{y}_{t,k+1})-f(\bm{y}_{t,k})]&\leq-\frac{\alpha}{2}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})\|^{2}]-\frac{\alpha}{2}\big{(}\frac{1}{2}-\alpha L\big{)}\mathbb{E}\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}+\frac{\alpha^{2}L}{2}\mathbb{E}\big{[}\|{\bm{d}}_{t,k}-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}\\ &\quad+\frac{\alpha}{2}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}\big{]}\\ &\quad+\underbrace{2\alpha\gamma^{2}L^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}\mathbb{E}\Big{[}\|\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\|^{2}\Big{]}}_{D_{t,k}}+\underbrace{\frac{2\alpha^{3}\beta^{2}L^{2}}{(1-\beta)^{2}}\mathbb{E}\big{[}\|\bm{m}_{t}\|^{2}\big{]}}_{M_{t}}.\end{split}$ (19)
We can now sum up (19) to bound progress within in one round of local updates. Next, we try to bound progress across outer iterations. From the definition of $\bm{y}_{t,k}$ in (12), we have
$\bm{y}_{t+1,0}-\bm{y}_{t,\tau}=\bm{x}_{t+1,0}-\bm{x}_{t,0}+\frac{\eta\gamma}{\tau R}\bm{m}_{t+1}=-\eta\gamma\mathcal{S}_{r}(\bm{m}_{t+1})+\frac{\eta\gamma}{\tau R}\bm{m}_{t+1}.$
By the $L$-smoothness of $f$, we have
$f(\bm{y}_{t+1,0})-f(\bm{y}_{t,\tau})\leq\langle\nabla f(\bm{y}_{t,\tau}),\bm{y}_{t+1,0}-\bm{y}_{t,\tau}\rangle+\frac{L\eta^{2}\gamma^{2}}{2}\|\frac{1}{\tau R}\bm{m}_{t+1}-\mathcal{S}_{r}(\bm{m}_{t+1})\|^{2}.$ (20)
Taking conditional expectation $\mathbb{E}_{t,\tau}$ (with randomness from $\mathcal{S}_{r}$) on (20), and using (14) and Lemma 1, we reach that
$\mathbb{E}_{t,\tau}\big{[}f(\bm{y}_{t+1,0})-f(\bm{y}_{t,\tau})\big{]}\leq\frac{dL\eta^{2}\gamma^{2}}{2}.$ (21)
Using (21), we are now prepared to bound the progress across outer iterations. Taking the total expectation and summing (19) and (21) from $k=0$ to $k=\tau-1$ and from $t=0$ to $t=T-1$, we obtain:
$\begin{split}\frac{\mathbb{E}[f(\bm{y}_{T,0})-f(\bm{y}_{0,0})]}{\tau T}&\leq-\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}-\frac{\alpha}{2\tau T}(\frac{1}{2}-\alpha L)\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}]\\ &\quad+\frac{1}{2}\alpha^{2}\zeta^{2}L+\frac{dL\eta^{2}\gamma^{2}}{2\tau}+\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]\\ &\quad+\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}D_{t,k}+\frac{1}{T}\sum_{t=0}^{T-1}M_{t}.\end{split}$ (22)
B.1.4 Bounding $D_{t,k}$
First, we have
$\|\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\|^{2}\leq 2\|\sum_{i=0}^{k-1}({\bm{d}}_{t,i}-\mathbb{E}_{t,i}[{\bm{d}}_{t,i}])\|^{2}+2k\sum_{i=0}^{k-1}\|\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}.$
Then, taking total expectation and summing over $k=0,\ldots,\tau-1$, from the above relation we have
$\begin{split}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\sum_{i=0}^{k-1}{\bm{d}}_{t,i}\|^{2}\big{]}&\leq 2\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\sum_{i=0}^{k-1}({\bm{d}}_{t,i}-\mathbb{E}_{t,i}[{\bm{d}}_{t,i}])\|^{2}\big{]}+2\sum_{k=0}^{\tau-1}k\sum_{i=0}^{k-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}\big{]}\\ &=2\sum_{k=0}^{\tau-1}\sum_{i=0}^{k-1}\mathbb{E}\big{[}\|{\bm{d}}_{t,i}-\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}\big{]}+2\sum_{k=0}^{\tau-1}k\sum_{i=0}^{k-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}\big{]}\\ &\leq 2\sum_{k=0}^{\tau-1}k\zeta^{2}+\tau(\tau-1)\sum_{i=0}^{k-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}\big{]}\\ &\leq\tau(\tau-1)\zeta^{2}+\tau(\tau-1)\sum_{i=0}^{k-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,i}[{\bm{d}}_{t,i}]\|^{2}\big{]}\end{split}$ (23)
where the first equality follows from the fact that the total expectation of crossing terms of $\|\sum_{i=0}^{k-1}({\bm{d}}_{t,i}-\mathbb{E}_{t,i}[{\bm{d}}_{t,i}])\|^{2}$ are zeros due to independence. From (23) we have
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}D_{t,k}\leq 2\alpha\tau(\tau-1)\gamma^{2}L^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}\Big{(}\frac{\zeta^{2}}{\tau}+\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}]\Big{)}.$ (24)
B.1.5 Bounding $M_{t}$
By the definition of $\bm{m}_{t}$ and $\bm{m}_{0}=\mathbf{0}$, we have
$\begin{split}\|\bm{m}_{t}\|^{2}&=\|\sum_{s=0}^{t-1}\beta^{t-1-s}(1-\beta)\Big{(}\sum_{k=0}^{\tau-1}{\bm{d}}_{s,k}\Big{)}\|^{2}\\ &\leq 2(1-\beta)^{2}\|\sum_{s=0}^{t-1}\beta^{t-1-s}\Big{(}\sum_{k=0}^{\tau-1}({\bm{d}}_{s,k}-\mathbb{E}_{s,k}[{\bm{d}}_{s,k}])\Big{)}\|^{2}+2(1-\beta)^{2}\|\sum_{s=0}^{t-1}\beta^{t-1-s}\Big{(}\sum_{k=0}^{\tau-1}\mathbb{E}_{s,k}[{\bm{d}}_{s,k}]\Big{)}\|^{2}.\end{split}$ (25)
We estimate the right two terms one by one. Taking total expectation and using the same arguments in equality of (23),
$\begin{split}&\mathbb{E}\Big{[}\|\sum_{s=0}^{t-1}\beta^{t-1-s}\Big{(}\sum_{k=0}^{\tau-1}({\bm{d}}_{s,k}-\mathbb{E}_{s,k}[{\bm{d}}_{s,k}])\Big{)}\|^{2}\Big{]}\\ =\ &\sum_{s=0}^{t-1}\beta^{2(t-1-s)}\mathbb{E}\Big{[}\|\sum_{k=0}^{\tau-1}({\bm{d}}_{s,k}-\mathbb{E}_{s,k}[{\bm{d}}_{s,k}]\|^{2}\Big{]}\\ =\ &\sum_{s=0}^{t-1}\beta^{2(t-1-s)}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|{\bm{d}}_{s,k}-\mathbb{E}_{s,k}[{\bm{d}}_{s,k}]\|^{2}\big{]}\\ \leq\ &\tau\zeta^{2}\sum_{s=0}^{t-1}\beta^{2(t-1-s)}\leq\frac{\tau\zeta^{2}}{1-\beta^{2}}.\end{split}$ (26)
Similarly, using Jensen’s inequality, we have
$\begin{split}&\mathbb{E}\Big{[}\|\sum_{s=0}^{t-1}\beta^{t-1-s}\Big{(}\sum_{k=0}^{\tau-1}\mathbb{E}_{s,k}[{\bm{d}}_{s,k}]\Big{)}\|^{2}\Big{]}\\ \leq\ &\tau\Big{(}\sum_{s=0}^{t-1}\beta^{t-1-s}\Big{)}\sum_{s=0}^{t-1}\beta^{t-1-s}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{s,k}{\bm{d}}_{s,k}\|^{2}\big{]}\\ \leq\ &\frac{\tau}{1-\beta}\sum_{s=0}^{t-1}\beta^{t-1-s}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{s,k}{\bm{d}}_{s,k}\|^{2}\big{]}.\end{split}$ (27)
Combing (25)-(LABEL:eq:mt-bound-p2) gives that
$\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}M_{t}&\leq\frac{4\tau\alpha^{3}\beta^{2}\zeta^{2}L^{2}}{1-\beta^{2}}+\frac{4\tau\alpha^{3}\beta^{2}L^{2}}{1-\beta}\frac{1}{T}\sum_{t=0}^{T-1}\sum_{s=0}^{t-1}\beta^{t-1-s}\sum_{k-0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{s,k}[{\bm{d}}_{s,k}]\|^{2}\big{]}\\ &=\frac{4\tau\alpha^{3}\beta^{2}\zeta^{2}L^{2}}{1-\beta^{2}}+\frac{4\tau\alpha^{3}\beta^{2}L^{2}}{1-\beta}\frac{1}{T}\sum_{t=0}^{T-2}\Big{(}\sum_{s=0}^{T-2-t}\beta^{s}\Big{)}\Big{(}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}\Big{)}\\ &\leq\frac{4\tau\alpha^{3}\beta^{2}\zeta^{2}L^{2}}{1-\beta^{2}}+\frac{4\tau\alpha^{3}\beta^{2}L^{2}}{(1-\beta)^{2}}\frac{1}{T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}.\end{split}$ (28)
B.1.6 Convergence rate
Putting (24) and (28) back into (22), we obtain that
$\begin{split}\frac{\mathbb{E}[f(\bm{y}_{T,0})-f(\bm{y}_{0,0})]}{\tau T}&\leq-\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}-\frac{\alpha}{2\tau T}(\frac{1}{2}-\alpha L)\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}]\\ &\quad+\frac{1}{2}\alpha^{2}\zeta^{2}L+\frac{dL\eta^{2}\gamma^{2}}{2\tau}+\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]\\ &\quad+2\alpha\tau(\tau-1)\gamma^{2}L^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}\Big{(}\frac{\zeta^{2}}{\tau}+\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}]\Big{)}\\ &\quad+\frac{4\tau\alpha^{3}\beta^{2}\zeta^{2}L^{2}}{1-\beta^{2}}+\frac{4\tau\alpha^{3}\beta^{2}L^{2}}{(1-\beta)^{2}}\frac{1}{T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}\\ &\leq-\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}-\frac{\alpha C_{1}}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\big{[}\|\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\big{]}\\ &\quad+\frac{\alpha}{2\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]+C_{2},\end{split}$ (29)
where
$C_{1}$ $=\frac{1}{4}-\frac{1}{2}\alpha L-2\tau(\tau-1)\gamma^{2}L^{2}\big{(}1-\frac{\eta}{\tau R}\big{)}^{2}-\frac{4\tau^{2}\alpha^{2}\beta^{2}L^{2}}{(1-\beta)^{2}},$
$C_{2}$ $=\alpha^{2}\zeta^{2}\Big{[}\frac{L}{2}+2\alpha(\tau-1)L^{2}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{4\tau\alpha\beta^{2}L^{2}}{1-\beta^{2}}\Big{]}+\frac{dL\eta^{2}\gamma^{2}}{2\tau}.$
When we take parameters to ensure that $C_{1}\geq 0$, from (29), we have
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}\leq\frac{2\mathbb{E}[f(\bm{y}_{0,0})-f(\bm{y}_{T,0})]}{\alpha\tau T}+\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]+\frac{2C_{2}}{\alpha}.$ (30)
Note that $\bm{y}_{0,0}=\bm{x}_{0,0}$ since $\bm{m}_{0}=\mathbf{0}$, and we denote $f_{*}=\inf_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})$. By taking $\alpha=\sqrt{\frac{n}{\tau T}}$ and recalling that $\alpha:=\frac{\eta\gamma}{\tau R}$, we obtain that
$\begin{split}\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}\leq\frac{2(f(\bm{x}_{0,0})-f_{*})}{\sqrt{n\tau T}}+\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})-\mathbb{E}[{\bm{d}}_{t,k}]\|^{2}]\\ +\zeta^{2}L\sqrt{\frac{n}{\tau T}}+\frac{4nL^{2}\zeta^{2}}{T}\Big{[}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{2\beta^{2}}{1-\beta^{2}}\Big{]}+dLR^{2}\sqrt{\frac{n\tau}{T}}.\end{split}$ (31)
Now we return to the constraint of $C_{1}$ when $\alpha=\sqrt{\frac{n}{\tau T}}$. Using $\alpha=\frac{\eta\gamma}{\tau R}$ in $C_{1}$ and rearranging terms gives that
$\frac{1}{4}-\frac{1}{2}\sqrt{\frac{nL^{2}}{\tau T}}-\frac{nL^{2}}{\tau T}\Big{[}2\tau(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{4\tau^{2}\beta^{2}}{(1-\beta)^{2}}\Big{]}\geq 0.$
The above relation reduces to
$\frac{\tau T}{nL^{2}}-2\sqrt{\frac{\tau T}{nL^{2}}}+1$ $\geq 8\tau(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{16\tau^{2}\beta^{2}}{(1-\beta)^{2}}+1,$
$\sqrt{\frac{\tau T}{nL^{2}}}$ $\geq\Big{[}8\tau(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{16\tau^{2}\beta^{2}}{(1-\beta)^{2}}+1\Big{]}^{1/2}+1.$
Since
$\Big{[}\Big{[}8\tau(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{16\tau^{2}\beta^{2}}{(1-\beta)^{2}}+1\Big{]}^{1/2}+1\Big{]}^{2}\leq 16\tau(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{32\tau^{2}\beta^{2}}{(1-\beta)^{2}}+4,$
it suffices to have
$T\geq 4nL^{2}\Big{[}4(\tau-1)\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{8\tau\beta^{2}}{(1-\beta)^{2}}+\frac{1}{\tau}\Big{]}.$
B.2 Proof of Theorem 2
Proof.
Since $\bm{d}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$, we have $\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(\bm{x}_{t,k}^{(i)})$. Then,
$\|\nabla f(\bm{x}_{t,k})-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}=\|\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(\bm{x}_{t,k})-\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(\bm{x}_{t,k}^{(i)})\|^{2}\leq\frac{L^{2}}{n}\sum_{i=1}^{n}\|\bm{x}_{t,k}-\bm{x}_{t,k}^{(i)}\|^{2},$ (32)
where in the last inequality we use Jensen’s inequality and $L$-smoothness of $f_{i}$’s. The quantity in (32) is a crucial in analyzing local SGD updates (Khaled et al., 2020). Since $\bm{x}_{t,0}^{(i)}$ is synchronized for every $t$, the local steps in Algorithm 1 is equivalent to that in local SGD. We can use the existing bound for deviations of local iterates in Wang et al. (2019) (equation (87)) or Yu et al. (2019) (Lemma 5 with $\beta=0$) in (32), then the term of effect of base optimizer in Theorem 1 can be bounded as
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\Big{[}\|\nabla f(\bm{x}_{t,k})-\mathbb{E}_{t,k}[{\bm{d}}_{t,k}]\|^{2}\Big{]}\leq\frac{L^{2}}{n\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\sum_{i=1}^{n}\mathbb{E}\big{[}\|\bm{x}_{t,k}-\bm{x}_{t,k}^{(i)}\|^{2}\big{]}\leq\frac{2\gamma^{2}L^{2}(\sigma^{2}\tau+3\delta^{2}\tau^{2})}{1-12\gamma^{2}L^{2}\tau^{2}}.$ (33)
When $\gamma L\tau\leq 1/6$, we have $21/(1-12\gamma^{2}L^{2}\tau^{2})\leq 3/$ 2. Combining this bound with (33), and using $\gamma=\frac{R}{\eta}\sqrt{\frac{n\tau}{T}}$, the bound in Theorem 1 reduces to
$\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}\leq\frac{2(f(\bm{x}_{0,0})-f_{*})}{\sqrt{n\tau T}}+\frac{3n\tau^{2}L^{2}R^{2}(\sigma^{2}+3\tau\delta^{2})}{\eta^{2}T}$
$+\sigma^{2}L\sqrt{\frac{n}{\tau T}}+\frac{4nL^{2}\sigma^{2}}{T}\Big{[}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{2\beta^{2}}{1-\beta^{2}}\Big{]}+dLR^{2}\sqrt{\frac{n\tau}{T}}.$
When $T\geq 36nL^{2}R^{2}\tau^{3}/\eta^{2}$, we have $\gamma L\tau\leq 1/6$, and thus $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}\|\nabla f(\bm{x}_{t,k})\|^{2}=O(1/\sqrt{T})$. ∎
B.3 Proof of Theorem 3
B.3.1 Preliminaries
Since we us SGD as the base optimizer, $\bm{d}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$. We define the following notations:
~$\tilde{\nabla}g(\bm{x}_{t,0})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)}),$ (34)
~$\tilde{\nabla}f(\bm{x}_{t,0})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{t,k}^{(i)}).$ (35)
We define the accumulative stochastic gradient noise at $t$-th outer iteration as:
$\bm{\delta}_{t}$ =~$=\tilde{\nabla}g(\bm{x}_{t,0})-\tilde{\nabla}f(\bm{x}_{t,0}).$
We use
$\bm{\epsilon}_{t+1}$ $=\bm{m}_{t+1}-\tau\nabla f(\bm{x}_{t,0}),$
to denote the gradient estimation error. Then, with $\gamma_{t}=\gamma$ for all $t$ an $\beta_{1}=\beta_{2}=\beta$, one has the following outer iteration for $t=0,\ldots,T-1$:
$\bm{m}_{t+1}$ $=\beta\bm{m}_{t}+(1-\beta)\tilde{\nabla}g(\bm{x}_{t,0}),$ (36)
$\bm{x}_{t+1,0}$ $=\bm{x}_{t,0}-\eta\gamma\mathrm{sign}(\bm{m}_{t+1}).$ (37)
We initialize
$\bm{m}_{0}=\mathbf{0}_{d},$
$\bm{x}_{0,0}=\bm{x}_{-1,0}-\eta\gamma\mathrm{sign}(\bm{m}_{0}).$
Hence $\bm{x}_{0,0}=\bm{x}_{-1,0}$ and $\bm{\epsilon}_{0}=-\tau\nabla f(\bm{x}_{t,0})$.
B.3.2 Descent Lemma
Lemma 2.
Under Assumption 1. For the iterates $\{\bm{x}_{t,0}\}$ in Algorithm 1 with $\beta_{1}=\beta_{2}=\beta,\lambda=0$, it satisfies that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{f(\bm{x}_{0,0})-f_{*}}{\eta\gamma T}+\frac{2\sqrt{d}}{\tau T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\bm{\epsilon}_{t+1}\|\big{]}+\frac{\eta\gamma dL}{2}.$ (38)
Proof.
The proof of this Lemma is adapted from Sun et al. (2023). From Assumption 1 we have
$f(\bm{x}_{t+1,0})$ $\leq f(\bm{x}_{t,0})+\nabla f(\bm{x}_{t,0})^{\top}(\bm{x}_{t+1,0}-\bm{x}_{t,0})+\frac{L}{2}\|\bm{x}_{t+1,0}-\bm{x}_{t,0}\|^{2}$
(37)${\leq}f(\bm{x}_{t,0})-\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}\mathrm{sign}(\bm{m}_{t+1})+\frac{\eta^{2}\gamma^{2}dL}{2}$
$\leq f(\bm{x}_{t,0})-\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}\mathrm{sign}(\nabla f(\bm{x}_{t,0}))+\eta\gamma\nabla f(\bm{x}_{t,0})^{\top}(\mathrm{sign}(\nabla f(\bm{x}_{t,0})-\mathrm{sign}(\bm{m}_{t+1}))+\frac{\eta^{2}\gamma^{2}dL}{2}$
$=f(\bm{x}_{t,0})-\eta\gamma\|\nabla f(\bm{x}_{t,0})\|_{1}+\eta\gamma\sum_{r=1}^{d}[\nabla f(\bm{x}_{t,0})]_{r}\cdot[\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})-\mathrm{sign}([\bm{m}_{t+1}]_{r})]+\frac{\eta^{2}\gamma^{2}dL}{2}.$
Recall that $\bm{\epsilon}_{t+1}=\bm{m}_{t+1}-\tau\nabla f(\bm{x}_{t,0})$. If $\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})\neq\mathrm{sign}([\bm{m}_{t+1}]_{r})$, then
$|[\bm{\epsilon}_{t+1}]_{r}|=|[\bm{m}_{t+1}]_{r}-\tau[\nabla f(\bm{x}_{t})]_{r}|\geq\tau|[\nabla f(\bm{x}_{t,0})]_{r}|,$
and thus
$[\nabla f(\bm{x}_{t,0})]_{r}\cdot[\mathrm{sign}([\nabla f(\bm{x}_{t,0})]_{r})-\mathrm{sign}([\bm{m}_{t+1}]_{r})]\leq\frac{2|[\bm{\epsilon}_{t+1}]_{r}|}{\tau}.$
It follows that
$f(\bm{x}_{t+1,0})\leq f(\bm{x}_{t,0})-\eta\gamma\|\nabla f(\bm{x}_{t,0})\|_{1}+\frac{2\eta\gamma\sqrt{d}\|\bm{\epsilon}_{t+1}\|}{\tau}+\frac{\eta^{2}\gamma^{2}dL}{2}.$
Summing over $t=0,\ldots,T-1$ and taking expectation, we obtain the desired relation in (38). ∎
B.3.3 Bounding estimation errors
We next try to bound the estimation error $\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]$. By the recursion of $\bm{m}_{t}$ and the definition of $\bm{\delta}_{t}$, for any $10\leq t\leq T-$ 1,
$\bm{m}_{t+1}$ $=\beta\bm{m}_{t}+(1-\beta)\tilde{\nabla}g(\bm{x}_{t,0})$
$=\beta(\bm{\epsilon}_{t}+\tau\nabla f(\bm{x}_{t-1,0}))+(1-\beta)(\bm{\delta}_{t}+\tau\nabla f(\bm{x}_{t,0}))+(1-\beta)(\tilde{\nabla}f(\bm{x}_{t,0})-\tau\nabla f(\bm{x}_{t,0}))$
Then, subtracting $\tau\nabla f(\bm{x}_{t,0})$ from both sides of the above relation gives that
$\bm{\epsilon}_{t+1}$ $=\beta\bm{\epsilon}_{t}+\beta\tau(\nabla f(\bm{x}_{t-1,0})-\nabla f(\bm{x}_{t,0}))+(1-\beta)\bm{\delta}_{t}+(1-\beta)(\tilde{\nabla}f(\bm{x}_{t,0})-\tau\nabla f(\bm{x}_{t,0}))$
$=\beta^{t+1}\bm{\epsilon}_{0}+\tau\sum_{s=0}^{t}\beta^{t+1-s}(\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0}))+(1-\beta)\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}$
$\quad+(1-\beta)\sum_{s=0}^{t}\beta^{t-s}(\tilde{\nabla}f(\bm{x}_{s,0})-\tau\nabla f(\bm{x}_{s,0})).$
Taking Euclidean norm and total expectation on the above relation leads to
$\begin{split}\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]&\leq\beta^{t+1}\mathbb{E}[\|\bm{\epsilon}_{0}\|]+\underbrace{\tau\sum_{s=0}^{t}\beta^{t+1-s}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0})\|\big{]}}_{E_{1}}+\underbrace{(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|\big{]}}_{E_{2}}\\ &\quad+\underbrace{(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}(\tilde{\nabla}f(\bm{x}_{s,0})-\tau\nabla f(\bm{x}_{s,0}))\|\big{]}}_{E_{3}}.\end{split}$ (39)
We next bound $E_{1},E_{2},E_{3}$ one by one.
Bounding $E_{1}$. From $L$-smoothness of $f$, we have, for all $s=0,\ldots,t$,
$\|\nabla f(\bm{x}_{s-1,0})-\nabla f(\bm{x}_{s,0})\|\leq L\|\bm{x}_{s-1,0}-\bm{x}_{s,0}\|\leq\eta\gamma\sqrt{d}L.$
Thus,
$E_{1}\leq\tau\eta\gamma\sqrt{d}L\sum_{s=0}^{t}\beta^{t+1-s}\leq\frac{\tau\eta\gamma\sqrt{d}L}{1-\beta}.$ (40)
Bounding $E_{2}$. By definition,
$E_{2}$ $=(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|\big{]}$
$\leq(1-\beta)\Big{(}\mathbb{E}[\|\sum_{s=0}^{t}\beta^{t-s}\bm{\delta}_{s}\|^{2}]\Big{)}^{1/2}$
$\leq(1-\beta)\Big{(}\mathbb{E}[\|\sum_{s=0}^{t}\beta^{t-s}(\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})-\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\nabla f_{i}(\bm{x}_{s,k}^{(i)}))\|^{2}]\Big{)}^{1/2}$
$\leq(1-\beta)\Big{(}\frac{1}{n^{2}}\sum_{s=0}^{t}\beta^{2t-2s}(\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})-\nabla f_{i}(\bm{x}_{s,k}^{(i)}))\|^{2}]\Big{)}^{1/2}$ (41)
$\leq\sqrt{\frac{\tau(1-\beta)}{n}}\sigma,$ (42)
where we used the mutual independence of stochastic gradient noise over workers, and the low of total expectations in (41), and used $\sqrt{1-\beta^{2}}>\sqrt{1-\beta}$ in (42).
Bounding $E_{3}$. By Assumption 1 and 3, we have
$\begin{split}E_{3}&=(1-\beta)\mathbb{E}\big{[}\|\sum_{s=0}^{t}\beta^{t-s}\big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\big{(}\nabla f_{i}(\bm{x}_{s,k}^{(i)})-\nabla f_{i}(\bm{x}_{s,0})\big{)}\big{]}\|\big{]}\\ &\leq(1-\beta)\frac{L}{n}\sum_{s=0}^{t}\beta^{t-s}\sum_{i=1}^{n}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\bm{x}_{s,k}^{(i)}-\bm{x}_{s,0}\|]\\ &\leq\frac{1}{2}(1-\beta)\tau(\tau-1)\eta LR\sum_{s=0}^{t}\beta^{t-s}\\ &\leq\frac{1}{2}\tau^{2}\eta LR,\end{split}$ (43)
where we used $\|\bm{x}_{s,k}^{(i)}-\bm{x}_{s,0}\|\leq\eta kR$ from the boundedness of $\|\nabla f_{i}(\bm{x}_{s,k}^{(i)},\bm{\xi}_{s,k}^{(i)})\|\leq R$. In light of (40)(42)(43), the relation (39) leads to
$\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]\leq\tau\beta^{t+1}\|\nabla f(\bm{x}_{0,0})\|+\frac{\tau\eta\gamma\sqrt{d}L}{1-\beta}+\sigma\sqrt{\frac{\tau(1-\beta)}{n}}+\frac{1}{2}\tau^{2}\eta LR.$ (44)
B.3.4 Proof of Theorem 3
We are now in a position to show Theorem 3.
Proof.
Using Lemma 2 with bound of $\mathbb{E}[\|\bm{\epsilon}_{t+1}\|]$ in (44), we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{f(\bm{x}_{0,0})-f_{*}}{\eta\gamma T}+\frac{2\sqrt{d}}{T(1-\beta)}\|\nabla f(\bm{x}_{0,0})\|+\frac{2dL\eta\gamma}{1-\beta}+2\sigma\sqrt{\frac{(1-\beta)d}{\tau n}}+\eta L\big{(}\tau R\sqrt{d}+\frac{\gamma d}{2}\big{)}.$
Taking $\eta=\frac{1}{LT^{3/4}}$, and $1-\beta=\frac{1}{\sqrt{T}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{L(f(\bm{x}_{0,0})-f_{*})}{\gamma T^{1/4}}+\frac{2\sqrt{d}}{T^{1/2}}\|\nabla f(\bm{x}_{0,0})\|+\frac{2d\gamma}{T^{1/4}}+\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}+\frac{\sqrt{d}\tau R+\gamma d/2}{T^{3/4}}.$
Appendix C Additional experiment details
C.1 Implementations
Our implementations [^1] https://github.com/shuhuayu/dist-sign-momentum build on codebases nanoGPT [^2] https://github.com/karpathy/nanoGPT, FairScale [^3] https://github.com/facebookresearch/fairscale, and Sophia [^4] https://github.com/Liuhong99/Sophia, where we use nanoGPT for GPT-2 (Radford et al., 2019) modeling, Sophia for the optimizer Sophia, and FairScale for distributed training modules. Our algorithm is implemented in the FairScale distributed training modules.
C.2 Additional experiment results
Comparison with local AdamW. In our experiments, we omit comparisons with Local AdamW, which uses AdamW as the base optimizer for local steps and periodically averages parameters across workers. In Figure 3, we compare the performance of Local AdamW with SlowMo and our proposed method under communication intervals of $\tau=12$ and $\tau=24$. The results show that Local AdamW is significantly slower than both SlowMo and our method, consistent with the findings of Sun et al. (2024).
Figure 3: Validation loss curves for communication interval $\tau=12,24.$
Comparisons of optimization errors. In addition to the improved generalization performance in terms of validation loss, we evaluate the optimization errors of our proposed method by plotting the training losses in Figure 4 for a communication interval of $\tau=12$. The results in Figure 4 demonstrate that our method consistently outperforms SlowMo in terms of optimization errors.
Figure 4: Training loss curves for communication interval $\tau=12$.
Additional validation loss curves. We plot the validation loss curves in Figures 5 for communication intervals $\tau=24$.
Figure 5: Validation loss curves for communication interval $\tau=24$.
Details of global AdamW step. The detailed description of a global update that mimics AdamW is presented in Algorithm 7.
Algorithm 7 Global AdamW with local steps. Component-wise vector multiplication $\bm{g}_{t}^{2}=\bm{g}_{t}\odot\bm{g}_{t}$, and $\sqrt{\widehat{\bm{v}}_{t}}$ means component-wise square root.
0:  Initialization $\bm{x}_{0,0}$; local update directions ${\bm{d}}_{t,k}^{(i)}$; local learning rate $\gamma$; global learning rate $\eta$; momentum coefficients $\beta_{1},\beta_{2}$; $\epsilon=10^{-8}$.
1:  $\bm{m}_{t}\leftarrow\mathbf{0},\bm{v}_{t}\leftarrow\mathbf{0}$
2:  for $t=0$ to $T-1$ do
3:     for $i=1$ to $n$ in parallel do
4:        for  $k=0$ to $\tau-1$ do
5:           $\bm{x}^{(i)}_{t,k+1}\leftarrow\bm{x}^{(i)}_{t,k}-\gamma_{t}{\bm{d}}_{t,k}^{(i)}$
6:        end for
7:     end for
8:     All-reduce step: $\bm{x}_{t,\tau}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}^{(i)}_{t,\tau}$
9:     Pseudo-gradient: $\bm{g}_{t}\leftarrow\frac{1}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$
10:     Momentum updates: $\qquad\bm{m}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+(1-\beta_{1})\bm{g}_{t}$ $\qquad\bm{v}_{t+1}\leftarrow\beta_{2}\bm{v}_{t}+(1-\beta_{2})\bm{g}_{t}^{2}$
11:     Bias-corrections: $\qquad\widehat{\bm{m}}_{t+1}\leftarrow\bm{m}_{t+1}/(1-\beta_{1}^{t+1})$ $\qquad\widehat{\bm{v}}_{t+1}\leftarrow\bm{v}_{t+1}/(1-\beta_{2}^{t+1})$
12:     Parameter update: $\qquad\bm{x}_{t+1,0}\leftarrow\bm{x}_{t,0}-\eta\Big{(}\frac{\widehat{\bm{m}}_{t+1}}{\sqrt{\widehat{\bm{v}}_{t+1}}+\epsilon}+\lambda\bm{x}_{t,0}\Big{)}$
13:     Synchronize over all workers: $\qquad\forall i\in[n],\ \bm{x}^{(i)}_{t+1,0}\leftarrow\bm{x}_{t+1,0}$
14:  end for
15:  return $\bm{x}_{T,0}$
On hyper-parameter fine-tuning. For AdamW, we fine-tune the peak LR and use the recommended values for $(\beta_{1},\beta_{2})$ and weight decay $\lambda$ from Liu et al. (2024b). Our results are consistent with those reported in Liu et al. (2024b). The same $(\beta_{1},\beta_{2})$ and $\lambda$ values are used for the base optimizer AdamW in both SlowMo and our proposed method. For SlowMo, we fine-tune the local peak LR, global LR, and momentum coefficients. In our proposed method, we adopt the values for $(\beta_{1},\beta_{2})$ and $\lambda$ suggested for Lion in Liu et al. (2024b), while fine-tuning only the global LR. Note that our proposed method is more sensitive to communication intervals due to its signed update nature, whereas in SlowMo, the momentum buffer can adapt to the accumulation of gradients from any number of local steps $\tau$, making it less sensitive to communication intervals.